MANIFOLDS NOT CONTAINING GOMPF NUCLEI

ANDRÁS I. STIPSICZ

Abstract. In this note we show that there are 4-manifolds not containing Gompf nucleus N_{2}; in this way we answer Problem 4.98 of Kirby’s problem list in the negative.